# Multi-UAV Behavior-based Formation with Static and Dynamic Obstacles Avoidance via Reinforcement Learning

### v1

@@ Abstract
Formation control of multiple Unmanned Aerial Vehicles (UAVs) is vital for practical applications. This paper tackles the task of behavior-based UAV formation while avoiding static and dynamic obstacles during directed flight. We present a two-stage reinforcement learning (RL) training pipeline to tackle the challenge of multi-objective optimization, large exploration spaces, and the sim-to-real gap. The first stage searches in a simplified scenario for a linear utility function that balances all task objectives simultaneously, whereas the second stage applies the utility function in complex scenarios, utilizing curriculum learning to navigate large exploration spaces. Additionally, we apply an attention-based observation encoder to enhance formation maintenance and manage varying obstacle quantity. Experiments in simulation and real world demonstrate that our method outperforms planning-based and RL-based baselines regarding collision-free rate and formation maintenance in scenarios with static, dynamic, and mixed obstacles.
@@ I Introduction
Figure 1: Illustration of the proposed two-stage training pipeline. In the first stage, we randomly search within the reward function space to find the best weight vector that balances all four objectives. The second stage applies the reward function to a more complex scenario and solves the transformed single-objective task with curriculum learning.
The formation control of multiple UAVs has received significant attention, showing great potential in practical applications such as search and rescue missions [1] and payload transport [2]. Prior studies on formation control can be sorted into three broad categories: leader-follower, virtual-structures, and behavior-based formation [3]. The leader-follower approach relies heavily on the leader; if the leader fails, the entire formation is compromised. The other two categories operate in a decentralized manner. In the virtual-structures method, UAVs follow individual reference targets to form a rigid formation, which restricts their flexibility. In contrast, behavior-based formation offers greater adaptability, since UAVs follow the formation’s centroid rather than individual reference points. Meanwhile, UAVs aim to maintain specified formation, avoid collisions with obstacles and other UAVs, and reach designated targets.
Compared with planning-based techniques[4, 5], multi-agent reinforcement learning (MARL) can produce more flexible behaviors, showing promise in solving difficult robotic tasks that requires aggressive maneuvers. Some research has successfully applied MARL to formation control with static obstacle avoidance in ground vehicles [6, 7]. Other studies use MARL for multi-UAV formation without considering obstacle avoidance [8, 9, 10]. Additionally, some work has employed RL to manage low-level actions, such as per-rotor thrusts, achieving robust obstacle avoidance without formation [11, 12]. In this paper, we tackle the more challenging task of maintaining behavior-based formation for multiple UAVs while avoiding both static and dynamic obstacles during directed flight via MARL. The primary challenges include:
  • Multi-objective optimization. The policy must balance several goals: achieving directed flight, maintaining behavior-based formation, avoiding static and dynamic obstacles, and preventing collisions with each other. These goals might conflict with each other in certain scenarios, making it challenging to design a reward function that incorporates the necessary trade-offs.
  • Large exploration space. The 3D nature of UAVs and task’s complexity considerably increase the exploration space.
  • Sim-to-real gap. This critical issue is magnified when RL generates low-level UAV instructions for high mobility to avoid dynamic obstacles.
@@
To address these challenges, we propose a novel two-stage RL training pipeline. In the first stage, we search for a linear utility function to balance four objectives, three for behavior completion and one for bridging the sim-to-real gap. In a simpler scenario, we evaluate the performance of different utility functions and select the one that best aligns with our preference. In the second stage, we apply the selected utility function to tackle a more complex task setting. Curriculum learning (CL), which progressively increases task difficulty, is utilized to accelerate training and enhance final performance. Additionally, we design an attention-based observation encoder to improve formation maintenance among multiple UAVs and to manage varying numbers of both dynamic and static obstacles.
Our experiments, carried out in both simulation and real-world environments, demonstrate that our method outperforms both planning-based and RL-based baselines regarding collision-free rate and formation maintenance when avoiding static, dynamic, and mixed obstacles. We also conduct ablation studies on CL and the attention-based observation encoder to validate the effectiveness of the design. Animated demonstrations can be viewed at our project website: https://sites.google.com/view/uav-formation-with-avoidance/.
@@ II-A Multi-UAV Formation
Extensive research has been conducted on multi-UAV formation. Classical planning-based methods [4, 5] and emerging learning-based methods [8, 9, 10, 13] guide UAVs to form a desired formation from random initial positions. Compared to the leader-follower scheme [14, 15], fully decentralized methods [16, 17, 18, 19] offer better scalability and resiliency to partial failures. However, these algorithms only consider obstacle-free or static-obstacle-only scenarios and do not account for sudden, dynamic obstacles such as flying birds. Additionally, some RL-based methods focus on navigating through static obstacles while maintaining formation for ground vehicles [6, 7]. In this paper, we derive an RL policy for behavior-based formation control in a fully decentralized manner. Transitioning from 2D to 3D space significantly increases the exploration space for RL algorithms, while extending from static to dynamic obstacles imposes higher demands on obstacle avoidance strategies.
@@ II-B Collision Avoidance for UAV
Obstacle avoidance is fundamental for UAVs in completing safe flights in cluttered environments. Methods like PANTHER [20] and Deep-PANTHER [21] effectively navigate a single UAV through obstacle-rich environments. In multi-UAV scenarios, UAVs must avoid various obstacles and potential collisions with other UAVs. Planning-based approaches utilize gradient-based optimization [22, 23, 24], evolutionary optimization [25], or particle swarm optimization [26] to plan collision-free paths for multi-UAVs in constrained environments like forests. However, these methods are designed for static environments, while real-world scenarios involve dynamic obstacles. Dynamic obstacle avoidance is particularly challenging, requiring UAVs to respond quickly to sudden obstacles and perform aggressive maneuvers [27, 28]. Planners typically need additional knowledge of obstacle trajectories, either given [29, 30] or predicted [31, 32, 33], for predictive planning. Although planning-based methods guarantee collision-free trajectories and can run in real-time, they constrain UAV behavior and may cause deadlocks. On the other hand, RL shows promise in overcoming these limitations for multi-UAV collision avoidance. [11] trains an RL policy to control motor thrust for multi-UAV tracking in obstacle-free environments, which can be adapted to collision avoidance by modeling obstacles as UAVs. Similarly, [12] treats multi-UAV collision avoidance as a single UAV obstacle avoidance problem. In this paper, we employ RL to tackle the complex task of multi-UAV collision avoidance in mixed obstacle scenarios while maintaining formation. We use multi-objective reinforcement learning to find the best weight between multiple objectives in simple scenarios and then apply the multi-objective weight to a more challenging setting, therefore balancing computation requirement and policy efficacy.
@@ III Task
Figure 2: Our task inherently contains four objectives: directed flight, behavior-based formation, obstacle avoidance, and action smoothing.
@@ III-A Task Setup
We address the task of maintaining behavior-based formation for multiple drones while avoiding both static and dynamic obstacles during directed flight. The task aims to achieve four objectives:
  • Directed Flight. The drones should fly together in a specified direction. Here we use target velocity as the command.
  • Behavior-based Formation. The drones should maintain a pre-set formation shape and size but can rotate the arrangement. For agile obstacle avoidance, the drones will not receive manually given target positions; rather, they have to determine their own positions based on the formation centroid and other drones’ status.
  • Obstacle Avoidance. The drones must avoid all environment obstacles, both dynamic and static, and prevent collisions with each other. The static obstacles are columns with radius of 0.15m and height of 3m. The dynamic obstacles are balls with radius of 0.15m, moving in a parabolic trajectory towards a random drone.
@@
Each objective corresponds to a reward function as detailed on the website. Together, they compose the reward vector $\textbf{R}=(R_{formation},R_{flight},R_{obstacle},R_{action})$.
@@ III-B Observation Space
The observation $\mathbf{o}^{i}$ for drone $i$ consists of the information of itself ($\mathbf{o}_{self}$), other drones ($\mathbf{o}_{drones}$), and obstacles ($\mathbf{o}_{static}$ and $\mathbf{o}_{dynamic}$). $\mathbf{o}_{self}$ consists of its position, orientation, linear and angular velocity, relative velocity compared to the reference velocity, and identifier. $\mathbf{o}_{drones}$ covers the L2 distance, relative position, and relative velocity of other drones relative to drone $i$. $\mathbf{o}_{static}$ uses a $3\times 3$ distance matrix to extract the relative position of the nearest obstacle, as proposed in [34]. $\mathbf{o}_{dynamic}$ includes the L2 distance, relative position, relative velocity, and absolute velocity of the obstacle.
@@ III-C Action Space
Inspired by [35], we adopt collective thrust and body rates (CTBR) commands as policy actions to ensure robust sim-to-real transfer and agile control. These CTBR commands are subsequently converted to motor thrust commands using PID controllers. Concretely, the action for drone $i$ is expressed as $\mathbf{a}=(c,\omega_{roll},\omega_{pitch},\omega_{yaw})$, where $c\in[0,1]$ indicates the collective thrust, and $\omega\in[-\pi,\pi]$ signifies the body rates for the corresponding axes.
@@ IV Preliminary
Solving the above MDP takes two steps. Firstly, we transform the multi-objective setting to a single-objective setting through a utility function $u:\mathbb{R}^{d_{obj}}\rightarrow\mathbb{R}$ and thus obtain a scalar reward $R$ function, $R:S\times A\rightarrow\mathbb{R}$. Then, as the above MDP becomes a standard Dec-POMDP, we apply multi-agent reinforcement learning algorithms to find an approximate solution.
@@ V Methodology
In this section, we detail our two-stage RL training pipeline that addresses the multi-objective task, as illustrated in Fig. 1. In the first stage, we identify a linear utility function that optimally aligns with our preference. Then, we transition to the second stage, where we use the utility function to tackle more complex task scenarios. We utilize curriculum learning (CL) to navigate large exploration spaces and accelerate training.
@@
We also describe the network architecture, which features an attention-based observation encoder to improve formation control and deal with different quantities of obstacles.
V-A Two-Stage Training Pipeline
V-A1 The First Stage, Reward Scalarization
Manually designing a reward function is difficult because the task involves multiple objectives with complex interrelationships. Therefore, we attempt to automatically transform the multi-objective reward into a scalar value. Specifically, we assume a linear utility function $u$, where the scalar reward $R$ is a weighted sum of the reward components, $=u(\textbf{R})=\textbf{w}^{\text{T}}\textbf{R}$ R , with weight vector $\textbf{w}=(w_{formation},w_{flight},w_{obstacle},w_{action})$ and reward vector $\textbf{R}=(R_{formation},R_{flight},R_{obstacle},R_{action})$. Using the weight vector w, we can establish the relationship between each objective, and thereby convert the multi-objective problem into a single-objective one.
@@ V-A1 The First Stage, Reward Scalarization
Existing multi-objective RL methods are not suitable for finding a proper w, as they cannot fully explore the weight space when the value of each objective becomes highly non-monotonic with the weight. Thus, we repeatedly select a random weight vector w, use the MAPPO algorithm [36] to solve the transformed Dec-POMDP, and evaluate the corresponding policy using satisfaction rate (SR) as the metric. An episode is satisfying if its four objectives meet certain thresholds, and SR is the proportion of satisfying episodes out of all testing episodes. We choose the w whose corresponding policy best aligns with our preference for the second stage. In practice, we select the w with the highest SR.
@@
Unfortunately, finding an appropriate weight vector requires multiple rounds of training, which is computationally expensive and time-consuming. To expedite the convergence of MAPPO for each w, we consider a simplified task: three drones maintaining a triangular formation while dodging three dynamic obstacles.
V-A2 The Second Stage, Reward Generalization

### v2

# Multi-UAV Formation Control with Static and Dynamic Obstacle Avoidance via Reinforcement Learning

@@ Abstract
This paper tackles the challenging task of maintaining formation among multiple unmanned aerial vehicles (UAVs) while avoiding both static and dynamic obstacles during directed flight. The complexity of the task arises from its multi-objective nature, the large exploration space, and the sim-to-real gap. To address these challenges, we propose a two-stage reinforcement learning (RL) pipeline. In the first stage, we randomly search for a reward function that balances key objectives: directed flight, obstacle avoidance, formation maintenance, and zero-shot policy deployment. The second stage applies this reward function to more complex scenarios and utilizes curriculum learning to accelerate policy training. Additionally, we incorporate an attention-based observation encoder to improve formation maintenance and adaptability to varying obstacle densities. Experimental results in both simulation and real-world environments demonstrate that our method outperforms both planning-based and RL-based baselines in terms of collision-free rates and formation maintenance across static, dynamic, and mixed obstacle scenarios. Ablation studies further confirm the effectiveness of our curriculum learning strategy and attention-based encoder. Animated demonstrations are available at: https://sites.google.com/view/uav-formation-with-avoidance/.
@@ I Introduction
Figure 1: Illustration of the proposed two-stage training pipeline. In the first stage, we randomly search within the reward function space to find the optimal function that balances all four objectives. In the second stage, we apply the selected reward function to a more complex scenario and solve the task with curriculum learning.
The collaboration of multiple unmanned aerial vehicles (UAVs) has received vast attention due to its significant potential in practical applications, such as search and rescue missions [1] and payload transportation [2]. However, this task is highly challenging, as it requires not only the basic capabilities of individual UAVs, such as flight control and obstacle avoidance, but also the effective coordination of all UAVs to achieve the collective objectives.
In this work, we focus on formation maintenance as the collective objective. Specifically, we address the challenge of multi-UAV formation control while ensuring both static and dynamic obstacle avoidance.
Obstacle avoidance is crucial in multi-UAV collaboration. In addition to navigating around environmental obstacles, as in single-UAV scenarios, UAVs in a multi-agent system must also avoid collisions with one another. Traditional planning-based methods [3, 4, 5, 6, 7] generate collision-free paths for UAVs to handle static obstacles. Dynamic obstacles, which are common in real-world environments, introduce additional complexity. Predictive planning for dynamic obstacles [8, 9, 10, 11, 12] often requires prior knowledge of obstacle trajectories, limiting its applicability. With the advancement of deep reinforcement learning (RL), learning-based methods [13, 14] have shown promising results in enabling multi-UAV collision avoidance in environments with both static and dynamic obstacles.
Formation control, the collective objective of our task, is also a critical aspect of multi-robot collaboration. However, when tackling formation control, the obstacle avoidance requirement is often simplified or overlooked. Planning-based methods [15, 16] typically focus on scenarios without obstacles or with only static obstacles. Similarly, learning-based approaches [17, 18, 19] utilize multi-agent reinforcement learning (MARL) for multi-UAV formation control in obstacle-free environments, and [20, 21] apply MARL to ground vehicle formation control with static obstacle avoidance only.
To the best of our knowledge, our work is the first to tackle the combined challenge of maintaining formation while avoiding both static and dynamic obstacles. This task is particularly challenging due to its multi-objective nature, where the UAVs must balance several potentially conflicting goals, such as directed flight, formation maintenance, and obstacle avoidance. Additionally, compared to ground vehicles, UAVs require their RL policy to tackle a larger sim-to-real gap and a broader exploration space, as they navigate in 3D space and have higher mobility.
@@
To address these challenges, we propose a two-stage reinforcement learning (RL) training pipeline. In the first stage, we search for a reward function that balances four task objectives, including directed flight, obstacle avoidance, formation maintenance, and action smoothness. The former three completes the task, and the last one bridges the sim-to-real gap. We evaluate the performance of different reward functions in a simpler scenario and select the one that best aligns with our preferences. In the second stage, we apply the selected reward function to tackle a more complex task setting. To accelerate training and enhance final performance, we employ curriculum learning (CL), which progressively increases task difficulty. Additionally, we design an attention-based observation encoder to handle varying numbers of obstacles and to improve formation maintenance among multiple UAVs.
Our main contributions are as follows:
  • We design a two-stage training pipeline to address the multi-objective task of formation maintenance and obstacle avoidance for both static and dynamic obstacles.
  • We conduct extensive simulations to demonstrate the effectiveness of our proposed training pipeline and network architecture. Our method outperforms both planning-based and RL-based baselines regarding collision-free rate and formation maintenance, and can easily generalize to varying number of obstacles.
  • We deploy the learned policy in real-world environment without additional fine-tuning, showcasing its practical applicability and robustness.
@@ II-A Multi-UAV Formation
Extensive research has been conducted on multi-UAV formation. Classical planning-based methods [15, 22] and emerging learning-based methods [17, 18, 19, 23] guide UAVs to form a desired configuration from random initial positions. Compared to the leader-follower schemes [24, 25], fully decentralized methods [26, 27, 28, 29] offer better scalability and resiliency to partial failures. However, these algorithms address only obstacle-free or static-obstacle scenarios and do not account for dynamic obstacles, such as flying birds. Additionally, some RL-based methods focus on navigating through static obstacles while maintaining formation in ground vehicle systems [20, 21]. In this paper, we derive an RL policy for UAV formation control in a fully decentralized manner. The transition from 2D to 3D space significantly enlarges the exploration space for RL algorithms, while incorporating dynamic obstacles introduces greater challenges for obstacle avoidance strategies.
@@ II-B Collision Avoidance for UAV
Obstacle avoidance is crucial for UAVs to ensure safe flights in cluttered environments. Methods like PANTHER [30] and Deep-PANTHER [31] effectively navigate a single UAV through obstacle-rich environments. In multi-UAV scenarios, UAVs must avoid not only environmental obstacles but also potential collisions with other UAVs. Planning-based approaches leverage gradient-based optimization [3, 4, 5], evolutionary optimization [6], or particle swarm optimization [7] to plan collision-free paths for multiple UAVs in constrained environments, such as forests. However, these methods are designed for static environments, while real-world scenarios involve dynamic obstacles. Avoiding dynamic obstacles is particularly challenging, as UAVs must respond quickly to sudden obstacles and execute aggressive maneuvers [32, 33]. Predictive planning methods usually require additional knowledge of obstacle trajectories, either provided [8, 9] or predicted [10, 11, 12]. Although planning-based approaches guarantee collision-free trajectories and can operate in real-time, they constrain UAV behavior and face a trade-off: decentralized planning may lead to deadlocks, while centralized planning reduces adaptability.
RL shows promise in overcoming these limitations for multi-UAV collision avoidance. [13] trains an RL policy to control motor thrust for multi-UAV tracking in obstacle-free environments, which can be adapted to collision avoidance by modeling obstacles as UAVs. Similarly, [14] treats multi-UAV collision avoidance as a single UAV obstacle avoidance problem. In this paper, we employ RL to tackle the complex task of multi-UAV collision avoidance in mixed obstacle scenarios while maintaining formation. We determine the optimal reward function for multiple objectives in simpler scenarios and then transfer this function to more challenging settings, thereby balancing computational requirements and policy efficacy.
@@ III-A Task Setup
We address the task of multi-UAV formation maintenance with obstacle avoidance during directed flight. The task aims to achieve four objectives:
  • Directed Flight. The drones should fly together in a specified direction, following a target velocity.
  • Formation Maintenance. The drones should maintain a predefined formation shape and size but can rotate the arrangement. For agile obstacle avoidance, the drones will not receive manually assigned target positions; instead, they must determine their own positions based on the formation centroid and the status of other drones.
  • Obstacle Avoidance. The drones must avoid all environment obstacles, both dynamic and static, and prevent collisions with each other.
@@
We initialize all drones at their respective positions within the desired formation. Multiple obstacles are placed along their flight path. We partition the environment using a zigzag grid arrangement, where each grid cell measures 0.5m with a 0.25m horizontal offset between adjacent rows, as shown in Fig. 1.
From these grids, we randomly sample multiple cells to place static obstacles, which are columns with a radius of 0.15m and height of 3m. Dynamic obstacles are balls with a radius of 0.15m, moving in a parabolic trajectory towards a randomly chosen drone.
@@ III-B Observation Space
The observation $\mathbf{o}^{i}$ for drone $i$ is composed of three components: self-information, $\mathbf{o}_{self}$; information about other drones, $\mathbf{o}_{drones}$; and details regarding static and dynamic obstacles, $\mathbf{o}_{static}$ and $\mathbf{o}_{dynamic}$.
$\mathbf{o}_{self}$ contains a 3D position coordinate $\mathbf{p}$, a 4D quaternion indicating its orientation $\mathbf{q}$, a 3D linear velocity $\mathbf{v}$, and a 3D angular velocity $\bm{\omega}$. To better capture the drone’s orientation, we rotate the x-axis and z-axis of the world coordinate using $\mathbf{q}$, producing 3D vectors $\mathbf{h}$ and $\mathbf{u}$. $\mathbf{o}_{self}$ also includes the drone’s relative velocity $\mathbf{v}_{rel}$ compared to the reference velocity $\mathbf{v}_{target}$ and a 3D representation of the drone’s identifier. $\mathbf{o}_{drones}$ covers the L2 distance, relative position, and relative velocity of other drones relative to drone $i$.
Following [34], we calculate the distance from $3\times 3$ grid cells around the drone to the nearest obstacles and use the $3\times 3$ distance tensor to characterize all obstacles. Unlike raw positional observations, this quantity- and permutation-invariant representation enhances the policy’s understanding of spatial data. We adopt this method and acquire a 9-dimensional $\mathbf{o}_{static}$ for each drone. $\mathbf{o}_{dynamic}$ includes the L2 distance, relative position, relative velocity, and absolute velocity of the obstacle.
@@ III-C Action Space
Inspired by [35], we adopt collective thrust and body rates (CTBR) commands as policy actions to ensure robust sim-to-real transfer and agile control. These CTBR commands are then converted to motor thrust commands using PID controllers. The action for drone $i$ is expressed as $\mathbf{a}=(c,\omega_{roll},\omega_{pitch},\omega_{yaw})$, where $c\in[0,0.9]*$max_thrust indicates the collective thrust, and $\omega\in[-\pi,\pi]$ signifies the body rates for the corresponding axes.
@@ IV Preliminary
Solving the above MDP takes two steps. Firstly, we transform the multi-objective setting to a single-objective setting through a utility function $u:\mathbb{R}^{d_{obj}}\rightarrow\mathbb{R}$ and thus obtain a scalar reward $R$ function, $R:S\times A\rightarrow\mathbb{R}$. Then, as the above MO-Dec-POMDP becomes a standard Dec-POMDP, we apply multi-agent reinforcement learning algorithms to find an approximate solution.
@@ V Methodology
We introduce a two-stage RL training pipeline to solve our proposed multi-objective task, as illustrated in Fig. 1. In the first stage, we identify a linear utility function $u(\textbf{R})=\textbf{w}^{\text{T}}\textbf{R}$ that properly balances between all objectives in a simplified scenario. Once the weight vector w is determined, we transition to the second stage, where we use the utility function to tackle more complex task scenarios. We use curriculum learning (CL) to navigate large exploration spaces and accelerate training.
@@
In this section, we detail the multi-objective reward function and our two-stage training pipeline. We also describe the network architecture, which features an attention-based observation encoder to improve formation control and deal with different quantities of obstacles.
V-A Multi-Objective Reward Function
As outlined in Sec. III-A, the task encompasses four objectives: directed flight, formation maintenance, obstacle avoidance, and action smoothing. Each objective is associated with a scalar reward component. Prior to detailing each reward component, we define three functions to characterize the relationship between reward $r$ and error $e$.
  • Linear: $r_{l}(e)=1-|e|$.
  • Reciprocal: $r_{r}(e)=1/(1+e)$.
  • Indicator: $r_{i}(e)=\mathbb{1}_{e>e_{thres}}$.
The first two functions ensure that when $e$ ranges from 0 to 1, $r$ remains within the same interval, and $r$ ascends as $e$ declines. The last function imposes a penalty when the error surpasses a specified threshold.
Flight Reward $R_{flight}$ involves constraints related to altitude error $e_{height}$, reference velocity error $e_{v}$, reference position error $e_{p}$, and heading error $e_{heading}$. The flight reward is expressed as $R_{flight}=\alpha_{heading}*r_{l}(e_{heading})+\alpha_{v}*r_{l}(e_{v})+\alpha_{p}*r_{r}(e_{p})+\alpha_{height}*r_{l}(e_{height})$.
Formation Reward $R_{formation}$ includes constraints related to formation size, shape, and the distances between drones. Following [26], we use the Laplacian distance $e_{shape}$ to measure formation similarity between the target formation and the current formation. We measure the pairwise distances between all drones, denoting the smallest as $dis$ and the largest as $size$. The size error is $e_{size}=(size-size_{target})^{2}$. Thus, we obtain $R_{formation}=\alpha_{shape}*r_{r}(e_{shape})+\alpha_{size}*r_{r}(e_{size})+\alpha_{dis}*r_{i}(dis)$.
Obstacle Reward $R_{obstacle}$ is determined by the distance from the drone to the closest obstacle. This reward only appears when obstacles are within the drone’s observation range.
Action Reward $R_{action}$ motivates the policy to generate smooth actions, quantified by the per-rotor difference in throttle between successive time steps. We evaluate the network output $e_{net}=\|\mathbf{a}_{t}-\mathbf{a}_{t-1}\|_{2}$, the drone’s throttle difference $e_{diff}=\|\mathbf{throttle}_{t}-\mathbf{throttle}_{t-1}\|_{2}$, the total throttle of all 4 rotors $e_{throt}=\sum\mathbf{throttle}_{t}$, and the rotation along the yaw axis $e_{yaw}$. Thus, the action reward is $R_{action}=\alpha_{net}*r_{l}(e_{net})+\alpha_{diff}*r_{l}(e_{diff})+\alpha_{throt}*r_{l}(e_{throt})+\alpha_{yaw}*r_{l}(e_{yaw})$.
Note that the $\alpha$ coefficients are set empirically to ensure that each term within the reward component is on a similar scale.
V-B Two-Stage Training Pipeline
V-B1 The First Stage, Reward Scalarization
In this stage, we transform the multi-objective vector reward $\textbf{R}=(R_{formation},R_{flight},R_{obstacle},R_{action})\in\mathbb{R}^{d_{obj}}$ into a scalar reward $R$ automatically. We assume $R$ is a weighted sum of the reward components, $=u(\textbf{R})=\textbf{w}^{\text{T}}\textbf{R}$ R, with weight vector $\textbf{w}=(w_{formation},w_{flight},w_{obstacle},w_{action})$, $\|\mathbf{w}\|_{1}=1$. The goal is thus to search for a set of w that best aligns with our preference.
@@ V-B1 The First Stage, Reward Scalarization
Existing MORL methods are not suitable for this problem as they cannot fully explore the weight space when the value of each objective becomes highly non-monotonic with the weight. Thus, to scalarize the reward vector, we repeatedly sample a random w and use MAPPO [36] to solve the transformed Dec-POMDP. For each w, we evaluate its corresponding RL policy using the satisfaction rate (SR) as the metric. An episode is considered satisfying if all four objectives meet predefined thresholds, and SR is the proportion of satisfying episodes out of the total episodes. In practice, we select the weight vector with the highest SR for the second stage.
@@
However, finding the appropriate weight vector requires multiple trials, which is computationally expensive and time-consuming. To accelerate the convergence of MAPPO in each trial, we simplify the task to three drones maintaining a triangular formation while avoiding three dynamic obstacles.
V-B2 The Second Stage, Reward Generalization